OpenStudy (anonymous):

Simplify the radical expression, I will give a medal to best answer and become a fan to anyone that helps!!

OpenStudy (anonymous):

\[(3+\sqrt{6)(3-\sqrt{6)}}\]

OpenStudy (anonymous):

a)-27 b)3 c)\[9+6\sqrt{6}\] d)\[15+6\sqrt{6}\]

OpenStudy (anonymous):

The answer is 3 :)

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (anonymous):

:D Welcome